# Simple homotopy types of even dimensional manifolds

### v2

@@ 4.2. Manifolds
By Corollary 3.9 (b), $\widehat{H}^{n+1}(C_{2};\operatorname{Wh}(G,w))=\operatorname{Im}\psi\subseteq\varrho^{-1}(\operatorname{Im}\sigma_{s})$. Hence $\varrho^{-1}(\operatorname{Im}\sigma_{s})=\widehat{H}^{n+1}(C_{2};\operatorname{Wh}(G,w))$, $(\varrho\circ\pi)^{-1}(\operatorname{Im}\sigma_{s})=\mathcal{J}_{n}(G,w)$. ∎
@@
Now we consider some corollaries of Theorems 4.11 and 4.16.
@@
$T(M)$ =$=\left\{\tau(g)\mid g\in\operatorname{hAut}(M)\right\}\subseteq\mathcal{J}_{n}(G,w)$
$U(M)$ =$=\left\{\pi(\tau(g))\mid g\in\operatorname{hAut}(M)\right\}\subseteq\widehat{H}^{n+1}(C_{2};\operatorname{Wh}(G,w))$
@@
Let $M$ be a closed, $\operatorname{CAT}$ $n$-manifold satisfying 1.1. Let $G=\pi_{1}(M)$ with orientation character $w\colon G\rightarrow\left\{\pm 1\right\}$. Then $|\mathcal{M}^{\operatorname{hCob}}_{s}(M)|>1$ if and only if $\mathcal{I}_{n}(G,w)\setminus T(M)$ is nonempty.
@@
Note that $T(M)=t_{M}(M)\subseteq\mathcal{J}_{n}(G,w)$ is the orbit of 00 under the action of $\operatorname{hAut}(M)$. So $q(\mathcal{I}_{n}(G,w))$ contains more than one element if and only if $\mathcal{I}_{n}(G,w)\setminus T(M)$ is nonempty. By Theorem 4.16 this is equivalent to $\mathcal{M}^{\operatorname{hCob}}_{s}(M)$ containing more than one element. ∎
@@
Let $M$ be a closed, $\operatorname{CAT}$ $n$-manifold satisfying 1.1. Let $G=\pi_{1}(M)$ with orientation character $w\colon G\rightarrow\left\{\pm 1\right\}$. Then $|\mathcal{M}^{h}_{s,\operatorname{hCob}}(M)|>1$ if and only if $\varrho^{-1}(\operatorname{Im}\sigma_{s})\setminus U(M)$ is nonempty. In particular, it is a sufficient condition that $\operatorname{Im}(\psi)\setminus U(M)$ is nonempty.
@@
Again $U(M)=u_{M}(M)\subseteq\widehat{H}^{n+1}(C_{2};\operatorname{Wh}(G,w))$ is the orbit of 00 under the action of $\operatorname{hAut}(M)$. So $\varrho^{-1}(\operatorname{Im}\sigma_{s})/\operatorname{hAut}(M)$ contains more than one element if and only if $\varrho^{-1}(\operatorname{Im}\sigma_{s})\setminus U(M)$ is nonempty. By Theorem 4.16 this is equivalent to $\mathcal{M}^{h}_{s,\operatorname{hCob}}(M)$ containing more than one element. Finally, by Corollary 3.9(b) we have $\varrho^{-1}(\operatorname{Im}\sigma_{s})\setminus U(M)\supseteq\operatorname{Im}(\psi)\setminus U(M)$. ∎
@@
Let $M$ be a closed, $\operatorname{CAT}$ $n$-manifold satisfying 1.1. Then the following are equivalent:
@@
    Either $|\mathcal{M}^{\operatorname{hCob}}_{s}(M)|>1$ or $|\mathcal{M}^{h}_{s,\operatorname{hCob}}(M)|>1$.
@@
    $(\varrho\circ\pi)^{-1}(\operatorname{Im}\sigma_{s})\setminus T(M)$ is nonempty.
@@
In this part we consider $S^{1}\times L$, the product of the circle with a lens space $L$. We will prove Theorems B, C, LABEL:, and D about the simple homotopy manifold sets of $S^{1}\times L$. As described in Sections 1.4 and 1.5, we need to study the involution on $\operatorname{Wh}(C_{\infty}\times C_{m})$ and the group $\operatorname{hAut}(S^{1}\times L)$. Section 5 contains our results about the groups $\mathcal{J}_{n}(C_{\infty}\times C_{m})$, $\mathcal{I}_{n}(C_{\infty}\times C_{m})$ and $\widehat{H}^{n+1}(C_{2};\operatorname{Wh}(C_{\infty}\times C_{m}))$ (for $n$ even), which rely on the results of Part III. In Section 6, we prove that all homotopy automorphisms of $S^{1}\times L$ are simple, and determine the automorphisms of $C_{\infty}\times C_{m}$ induced by them. In Section 7, we combine these results to prove Theorems B, C, LABEL:, and D.
@@ 5. The involution on $\operatorname{Wh}(C_{\infty}\times C_{m})$
We will use the version given by Ranicki [Ra86]. We start by defining the terms appearing in the decomposition in Section 5.1. Section 5.2 contains the decomposition of $\operatorname{Wh}(C_{\infty}\times C_{m})$ and the induced decompositions of $\mathcal{J}_{n}(C_{\infty}\times C_{m})$, $\mathcal{I}_{n}(C_{\infty}\times C_{m})$ and $\widehat{H}^{n+1}(C_{2};\operatorname{Wh}(C_{\infty}\times C_{m}))$. These will be combined with the results of Part III in Section 5.3, to prove Theorems 5.14, 5.15, LABEL:, and 5.17, which will be the key algebraic ingredients in the proofs of Theorems B, C, LABEL:, and D respectively.
@@ 5.1. The $\widetilde{K}_{0}$ and $NK_{1}$ groups
First we consider the $K_{0}$ groups. For a ring $R$, we will define the abelian group $K_{0}(R)$, and, for a group $G$, the reduced group $\widetilde{K}_{0}(\mathbb{Z}G)$. A convenient reference for this material is [We13]. In Part III, we will study the involution on $\widetilde{K}_{0}(\mathbb{Z}G)$ in more detail when $G$ is a finite cyclic group.
@@
For a ring $R$ let $P(R)$ denote the set of isomorphism classes of (finitely generated left) projective $R$-modules, which is a monoid under direct sum. Define $K_{0}(R)$ to be the Grothendieck group of this monoid, i.e. the abelian group generated by symbols $[P]$, for every $P\in P(R)$, subject to the relations $[P_{1}\oplus P_{2}]=[P_{1}]+[P_{2}]$ for $P_{1},P_{2}\in P(R)$.
@@
For a group $G$, let
$\widetilde{K}_{0}(\mathbb{Z}G)\coloneqq K_{0}(\mathbb{Z}G)/K_{0}(\mathbb{Z}),$
where the map $K_{0}(\mathbb{Z})\to K_{0}(\mathbb{Z}G)$ is induced by the inclusion $i\colon\mathbb{Z}\rightarrow\mathbb{Z}G$.
@@
For a group $G$ and a (finitely generated) projective $\mathbb{Z}G$-module $P$, the rank of $P$ is defined to be
@@
where $\varepsilon\colon\mathbb{Z}G\to\mathbb{Z}$ denotes the augmentation map and $\operatorname{rk}_{\mathbb{Z}}$ denotes the rank of a free $\mathbb{Z}$-module.
We have that $K_{0}(\mathbb{Z})\cong\mathbb{Z}$ and, for $P\in P(\mathbb{Z}G)$, the rank $\operatorname{rk}(P)$ coincides the the image of $[P]$ under the composition $K_{0}(\mathbb{Z}G)\to K_{0}(\mathbb{Z})\cong\mathbb{Z}$ induced by $\varepsilon$. Since $\varepsilon\circ i=\operatorname{Id}_{\mathbb{Z}}$, we have a splitting of abelian groups $K_{0}(\mathbb{Z}G)\cong\mathbb{Z}\oplus\widetilde{K}_{0}(\mathbb{Z}G)$ given by $[P]\mapsto(\operatorname{rk}(P),[P])$. In particular, $\widetilde{K}_{0}(\mathbb{Z}G)$ is the set of equivalence classes of projective $\mathbb{Z}G$-modules $P$ where $P\sim Q$ if $P\oplus\mathbb{Z}G^{i}\cong Q\oplus\mathbb{Z}G^{j}$ for some $i,j\geq 0$.
@@
The standard involution on $K_{0}(R)$ is given by $[P]\mapsto-[P^{*}]$. This map preserves the rank of projective $\mathbb{Z}G$-modules and so induces an involution on $\widetilde{K}_{0}(\mathbb{Z}G)$.
@@
$NK_{1}(R)\coloneqq\operatorname{coker}(K_{1}(R)\rightarrow K_{1}(R[t])),$
@@
Note that $NK_{1}$ is a functor from the category of rings to the category of abelian groups. If $R$ (and hence $R[t]$) is a ring with involution, then $K_{1}(R[t])$ also has a natural involution (see Section 2.2). This involution preserves the image of $K_{1}(R)\rightarrow K_{1}(R[t])$, and hence induces an involution on $NK_{1}(R)$, which we will denote by $x\mapsto\overline{x}$.
@@
We will need the following result of Farrell [Fa77] (note that $\mathrm{Nil}$, which is also written as $\mathrm{Nil}_{0}$, coincides with $NK_{1}$ by [Ba68, Chapter XII; 7.4(a)]).
@@
The following can be proven by combining results of Bass-Murthy [BM67], Martin [Ma75], and Weibel [We09].

### v3

@@ 4.2. Manifolds
By Corollary 3.10, we have $\widehat{H}^{n+1}(C_{2};\operatorname{Wh}(G,w))=\operatorname{Im}\psi\subseteq\varrho^{-1}(\operatorname{Im}\sigma_{\mathrm{s}})$. Hence $\varrho^{-1}(\operatorname{Im}\sigma_{\mathrm{s}})=\widehat{H}^{n+1}(C_{2};\operatorname{Wh}(G,w))$ and $(\varrho\circ\pi)^{-1}(\operatorname{Im}\sigma_{\mathrm{s}})=\mathcal{J}_{n}(G,w)$. ∎
@@
We now consider some corollaries of Theorems 4.11 and 4.16.
@@
$T(M)$ $=\left\{\tau(g)\mid g\in\operatorname{hAut}(M)\right\}\subseteq\mathcal{J}_{n}(G,w)$
$U(M)$ $=\left\{\pi(\tau(g))\mid g\in\operatorname{hAut}(M)\right\}\subseteq\widehat{H}^{n+1}(C_{2};\operatorname{Wh}(G,w))$
@@
Let $M$ be a closed, $\operatorname{CAT}$ $n$-manifold with $\operatorname{CAT}$ as in 1.1. Let $G=\pi_{1}(M)$ with orientation character $w\colon G\rightarrow\left\{\pm 1\right\}$. Then $|\mathcal{M}^{\operatorname{hCob}}_{\mathrm{s}}(M)|>1$ if and only if $\mathcal{I}_{n}(G,w)\setminus T(M)$ is nonempty.
@@
Note that $T(M)=t_{M}(M)\subseteq\mathcal{J}_{n}(G,w)$ is the orbit of 0 under the action of $\operatorname{hAut}(M)$. So $q(\mathcal{I}_{n}(G,w))$ contains more than one element if and only if $\mathcal{I}_{n}(G,w)\setminus T(M)$ is nonempty. By Theorem 4.16 this is equivalent to $\mathcal{M}^{\operatorname{hCob}}_{\mathrm{s}}(M)$ containing more than one element. ∎
@@
Let $M$ be a closed, $\operatorname{CAT}$ $n$-manifold with $\operatorname{CAT}$ as in 1.1. Let $G=\pi_{1}(M)$ with orientation character $w\colon G\rightarrow\left\{\pm 1\right\}$. Then $|\mathcal{M}^{\mathrm{h}}_{\mathrm{s},\operatorname{hCob}}(M)|>1$ if and only if $\varrho^{-1}(\operatorname{Im}\sigma_{\mathrm{s}})\setminus U(M)$ is nonempty. In particular, it is a sufficient condition that $\operatorname{Im}(\psi)\setminus U(M)$ is nonempty.
@@
Again $U(M)=u_{M}(M)\subseteq\widehat{H}^{n+1}(C_{2};\operatorname{Wh}(G,w))$ is the orbit of 0 under the action of $\operatorname{hAut}(M)$. So $\varrho^{-1}(\operatorname{Im}\sigma_{\mathrm{s}})/\operatorname{hAut}(M)$ contains more than one element if and only if $\varrho^{-1}(\operatorname{Im}\sigma_{\mathrm{s}})\setminus U(M)$ is nonempty. By Theorem 4.16 this is equivalent to $\mathcal{M}^{\mathrm{h}}_{\mathrm{s},\operatorname{hCob}}(M)$ containing more than one element. Finally, by Corollary 3.10 we have $\varrho^{-1}(\operatorname{Im}\sigma_{\mathrm{s}})\setminus U(M)\supseteq\operatorname{Im}(\psi)\setminus U(M)$. ∎
@@
Let $M$ be a closed, $\operatorname{CAT}$ $n$-manifold with $\operatorname{CAT}$ as in 1.1. Then the following are equivalent:
@@
    Either $|\mathcal{M}^{\operatorname{hCob}}_{\mathrm{s}}(M)|>1$ or $|\mathcal{M}^{\mathrm{h}}_{\mathrm{s},\operatorname{hCob}}(M)|>1$.
@@
    $(\varrho\circ\pi)^{-1}(\operatorname{Im}\sigma_{\mathrm{s}})\setminus T(M)$ is nonempty.
@@
Remark 4.23.
We can relate simple homotopy manifold sets to the structure sets discussed in Section 3. As $\mathcal{S}^{\mathrm{h}}(M)$ contains all homotopy equivalences $f\colon N\rightarrow M$ up to $h$-cobordism, the forgetful map $f]\mapsto[N]$ is a surjection $\mathcal{S}^{\mathrm{h}}(M)\rightarrow\mathcal{M}^{\mathrm{h}}_{\mathrm{s},\operatorname{hCob}}(M)$. Moreover, $\widehat{\tau}\colon\mathcal{S}^{\mathrm{h}}(M)\to\widehat{H}^{n+1}(C_{2};\operatorname{Wh}(G,w))$ descends to the map $\widetilde{u}_{M}$ (cf. the proof of Theorem 4.11 (e)). Similarly, there are surjections $\mathcal{S}^{\mathrm{hCob}}_{\mathrm{sCob}}(M)\rightarrow\mathcal{M}^{\operatorname{hCob}}_{\mathrm{s}}(M)$ and $\mathcal{S}^{\mathrm{h}}_{\mathrm{sCob}}(M)\rightarrow\mathcal{M}^{\mathrm{h}}_{\mathrm{s}}(M)$ (see Remark 3.12 for the definitions of these structure sets), and $\widetilde{t}^{\prime}_{M}$ and $\widetilde{t}_{M}$ are induced by the maps $\tau$ that take the Whitehead torsion of a homotopy equivalence (induced by the $h$-cobordism in the case of $\mathcal{S}^{\mathrm{hCob}}_{\mathrm{sCob}}(M)$). These observations, and the realisation maps from Remark 3.12, are summarised in the following commutative diagram.
${\operatorname{Wh}(G,w)\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\begin{subarray}{c}x\mapsto\\ -x+\\ (-1)^{n}\overline{x}\end{subarray}}$$\scriptstyle{W_{\operatorname{Wh}}}$${L_{n+1}^{\mathrm{s},\tau}(\mathbb{Z}G,w)\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{W_{\mathrm{s},\tau}}$${L_{n+1}^{\mathrm{h}}(\mathbb{Z}G,w)\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\psi}$$\scriptstyle{W_{\mathrm{h}}}$${\mathcal{S}^{\mathrm{hCob}}_{\mathrm{sCob}}(M)\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$  $\scriptstyle{\tau}$  ${\mathcal{S}^{\mathrm{h}}_{\mathrm{sCob}}(M)\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$  $\scriptstyle{\tau}$  ${\mathcal{S}^{\mathrm{h}}(M)\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\widehat{\tau}}$${\mathcal{I}_{n}(G,w)\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{q}$${\mathcal{J}_{n}(G,w)\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\pi}$$\scriptstyle{q}$${\widehat{H}^{n+1}(C_{2};\operatorname{Wh}(G,w))\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${\mathcal{M}^{\operatorname{hCob}}_{\mathrm{s}}(M)\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$  $\scriptstyle{\widetilde{t}^{\prime}_{M}}$${\mathcal{M}^{\mathrm{h}}_{\mathrm{s}}(M)\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$  $\scriptstyle{\widetilde{t}_{M}}$${\mathcal{M}^{\mathrm{h}}_{\mathrm{s},\operatorname{hCob}}(M)\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\widetilde{u}_{M}}$${q(\mathcal{I}_{n}(G,w))\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${\mathcal{J}_{n}(G,w)/\operatorname{hAut}(M)\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${\widehat{H}^{n+1}(C_{2};\operatorname{Wh}(G,w))/\operatorname{hAut}(M)}$
@@
In this part we consider $S^{1}\times L$, the product of the circle with a lens space $L$. We prove Theorems C, D, LABEL:, and E about the simple homotopy manifold sets of $S^{1}\times L$. As described in Theorems B and 1.4, we need to study the involution on $\operatorname{Wh}(C_{\infty}\times C_{m})$ and the group $\operatorname{hAut}(S^{1}\times L)$. Section 5 contains our results about the groups $\mathcal{J}_{n}(C_{\infty}\times C_{m})$, $\mathcal{I}_{n}(C_{\infty}\times C_{m})$ and $\widehat{H}^{n+1}(C_{2};\operatorname{Wh}(C_{\infty}\times C_{m}))$ (for $n$ even), which rely on the results of Part III. In Section 6, we prove that all homotopy automorphisms of $S^{1}\times L$ are simple, and determine the automorphisms of $C_{\infty}\times C_{m}$ induced by them. In Section 7, we combine these results to prove Theorems C, D, LABEL:, and E.
@@ 5. The involution on $\operatorname{Wh}(C_{\infty}\times C_{m})$
We use the version given by Ranicki [Ra86]. We start by defining the terms appearing in the decomposition in Section 5.1. Section 5.2 contains the decomposition of $\operatorname{Wh}(C_{\infty}\times C_{m})$ and the induced decompositions of $\mathcal{J}_{n}(C_{\infty}\times C_{m})$, $\mathcal{I}_{n}(C_{\infty}\times C_{m})$ and $\widehat{H}^{n+1}(C_{2};\operatorname{Wh}(C_{\infty}\times C_{m}))$. These are combined with the results of Part III in Section 5.3, to prove Theorems 5.14 and 5.15, which are the key algebraic ingredients in the proofs of Theorems C, D, LABEL:, and E.
@@ 5.1. The $\widetilde{K}_{0}$ and $NK_{1}$ groups
First we consider the $K_{0}$ groups. For a ring $R$, we define the abelian group $K_{0}(R)$ and, for a group $G$, the reduced group $\widetilde{K}_{0}(\mathbb{Z}G)$. A convenient reference for this material is [We13]. In Part III, we study the involution on $\widetilde{K}_{0}(\mathbb{Z}G)$ in more detail when $G$ is a finite cyclic group.
@@
For a ring $R$ let $P(R)$ denote the set of isomorphism classes of finitely generated left projective $R$-modules, which is a monoid under direct sum. Define $K_{0}(R)$ to be the Grothendieck group of this monoid, i.e. the abelian group generated by symbols $[P]$, for every $P\in P(R)$, subject to the relations $P_{1}\oplus P_{2}]=[P_{1}]+[P_{2}]$ for $P_{1},P_{2}\in P(R)$.
@@
We say that a ring $R$ is $\mathbb{Z}$-augmented if it comes equipped with a ring homomorphism $\varepsilon\colon R\to\mathbb{Z}$. If $i\colon\mathbb{Z}\to R$ is the unique ring homomorphism, then $\varepsilon\circ i=\operatorname{Id}_{\mathbb{Z}}$ and so $i$ is injective and $\varepsilon$ is surjective. For example, $\mathbb{Z}G$ is $\mathbb{Z}$-augmented where $\varepsilon\colon\mathbb{Z}G\to\mathbb{Z}$ denotes the augmentation map.
@@
For a $\mathbb{Z}$-augmented ring $R$, let
$\widetilde{K}_{0}(R):=K_{0}(R)/K_{0}(\mathbb{Z}),$
where the map $K_{0}(\mathbb{Z})\to K_{0}(R)$ is induced by the inclusion $i\colon\mathbb{Z}\rightarrow R$.
@@
For a $\mathbb{Z}$-augmented ring $R$ and a finitely generated projective $R$-module $P$, the rank of $P$ is defined to be
@@
where $\varepsilon\colon R\to\mathbb{Z}$ denotes the augmentation map and $\operatorname{rk}_{\mathbb{Z}}$ denotes the rank of a free $\mathbb{Z}$-module.
Let $R$ be a $\mathbb{Z}$-augmented ring. We have that $K_{0}(\mathbb{Z})\cong\mathbb{Z}$ and, for $P\in P(R)$, the rank $\operatorname{rk}(P)$ coincides with the image of $[P]$ under the composition $K_{0}(R)\to K_{0}(\mathbb{Z})\cong\mathbb{Z}$ induced by $\varepsilon$. Since $\varepsilon\circ i=\operatorname{Id}_{\mathbb{Z}}$, we have a splitting of abelian groups $K_{0}(R)\cong\mathbb{Z}\oplus\widetilde{K}_{0}(R)$ given by $P]\mapsto(\operatorname{rk}(P),[P])$. In particular, $\widetilde{K}_{0}(R)$ is the set of equivalence classes of finitely generated projective $R$-modules $P$ where $P\sim Q$ if $P\oplus R^{i}\cong Q\oplus R^{j}$ for some $i,j\geq 0$.
@@
The standard involution on $K_{0}(R)$ is given by $P]\mapsto-[P^{*}]$. If $R$ is a $\mathbb{Z}$-augmented ring, this map preserves the rank of a projective $R$-module and so induces an involution on $\widetilde{K}_{0}(R)$.
@@
$NK_{1}(R):=\operatorname{coker}(K_{1}(R)\rightarrow K_{1}(R[t])),$
@@
Note that $NK_{1}$ is a functor from the category of rings to the category of abelian groups. If $R$ is a ring with involution, then $R[t]$ has an involution such that $t\mapsto t^{-1}$ and which extends the involution on $R$. Thus $K_{1}(R[t])$ has a natural involution (see Section 2.2), and this involution preserves the image of $K_{1}(R)\rightarrow K_{1}(R[t])$, and hence induces an involution on $NK_{1}(R)$, which we denote by $x\mapsto\overline{x}$.
@@
We need the following result of Farrell [Fa77] (note that $\mathrm{Nil}$, which is also written as $\mathrm{Nil}_{0}$, coincides with $NK_{1}$ by [Ba68, Chapter XII; 7.4(a)]).
@@
We conclude this section with the following result.